# Labelling Data with Unknown References

### v2

@@ 7.1 Correctness
Remark 3.
The exponential bound from Lemma 1 means that, for example, after running the EV protocol for 3 rounds, the probability that the verifier fails to detect a lie from the evaluator is 1.6%.
@@
The implication of Theorem 1 is that a good evaluator with low probability to flip the labels will obtain a good label accuracy. A bad evaluator with high probability to flip the labels will also obtain a good accuracy. Another interpretation of this, however, is that there is no way to recover unknowable labels when the evaluator is unable to convince the verifier–we discuss this further in Section 9.
It is worth mentioning that the assumption for Theorem 1 around always lying when mislabelling (and hence always being wrong) are not necessarily applicable to all evaluators, since a particularly tricky evaluator could still guess the right label and yet not pass the challenges. We consider this out of scope, since the ability to guess a label without explaining how is not a viable way to establish trust in scenarios without any references. Consequentially, for the No-Data Algorithm it is more helpful to look at the success count.
@@ 8 Experiments
We describe the two experiments and results in this section and refer the reader to Appendix D for a more detailed treatment, including prompts and specifics on the models used. Throughout this section, we run the No-Data Algorithm with $r=3$.
8.0.1 Data
We created two mostly disjoint synthetic datasets so that we could control the labels, data, and their rubrics. We refer to these datasets as in-phenomenon, or IP; and out-of-phenomenon, or OOP, as opposed to in-distribution (r. out-of-distribution) to highlight the fact that the entire measurement is altered by the change in data–and, that, semantically, we are modelling different phenomena. The IP dataset is our baseline: the evaluator will not be lying on knowing this phenomenon. The OOP dataset is our reference for a corpus that the evaluator has not seen before (and thus it will lie about knowing it). For both problems, $\sigma$ is the majority vote.
We chose to generate our own data to ablate out any potential memorisation concerns, and to control the rubrics. These may be found in Table 1. The test sets for both IP and OOP included 498 entries and are balanced.
8.0.2 Evaluator
We used a decision tree (DT) for our evaluation, and an LLM (o3-mini; [31]). Both models observed a balanced training split of the IP dataset: the DT was trained on it, while the LLM observed five exemplars in the prompt. The rationale to train the DT was to fully ensure that the evaluator learnt both $\sigma$ and $C$. For generation it only had access to the rubric, but not $\sigma$. The LLM observed it via the prompt, and was made aware of the aggregation function only for the labelling step. For generation it was asked to pick the best candidate from the dataset, as opposed to outputting a new point, as we observed marked performance decreases otherwise. During OOP, the LLM was given the rubric for IP and an OOP datapoint; while the DT was trained in IP (and, ideally, learnt the rubric), and classified IP and OOP datapoints.
@@ 8.0.2 Evaluator
$c_{0}$ $y=1$ if $x$ has an even number of ones; else 00
$c_{1}$ $y=1$ if $x$ starts with a zero or contains $10101$, but not both; else 00
$c_{2}$ $y=1$ if $x$ has strictly more than five ones; else 00
@@
$c_{0}$ $y=1$ if $x$ contains the substring $111$; else 00
$c_{1}$ $y=1$ if $x$ ends with a one; else 00
$c_{2}$ $y=1$ if $x$ contains the substring $110001$; else 00
Table 1: Rubrics and criteria for our experiments. For both problems, $\sigma$ was the majority vote. Remark that $c_{1}$ for the IP rubric is an exclusive-or operation. This means that, while this rubric contains three criteria, the total rubric contains five: $\{c_{0},c_{2},c_{1,a},c_{1,b},c_{1}\}$; where $c_{1,a},c_{1,b}$ are the clauses for $c_{1}$, and $c_{1}$ is the exclusive-or itself.
8.0.3 Verifier
The verifier is a rule-matching algorithm that has access to the relevant rubrics (e.g., when testing IP, it observes the rubric for IP), and works as described in Section 6.1.
8.1 Results
The results are in Table 2. Both the DT and the LLM adjusted to the predictions correctly, successfully convincing the verifier when it was an oracular generator, reaching approximately their knowable accuracy, and outputting low success rates when lying. Indeed, the accuracy of the models in the OOP dataset hovers around 54%–close to random guessing. A similar accuracy can be obtained by the output of the No-Data Algorithm, but the success rates between IP (100%) and OOP (5%) provide better information as to in which situations the evaluator is trustworthy. The tuned $\phi$ value, based on the known labels, also yielded accuracies close to the expected score from the known dataset. Of note, which we will cover in Section 9, is the number of flips performed by the algorithm in the DT OOP dataset: 46%.
As part of our experimentation, we performed ablation studies on the relationship of lying with evaluators and generators (Section 8.1.1); and the effectiveness of generator strategies in LLMs (Section 8.1.2).
8.1.1 Ablation: Lying Evaluators and Lying Generators
  1. 1.
  2. 2.
  3. 3.
    approximately know $f$ up to a probability $p=1/10$ (r. it could pass both challenges with probability $1-p$).
As before, the results indicate that lying evaluators tend to score low in the success rate, even when their accuracies are close to the known case. Remark that (a) the evaluator that lies on knowing $\sigma$ had a success rate much higher than that of lying of knowing $f$; and (b) the evaluator with approximately knowledge of $f$ had a comparatively low (44%) success rate. We discuss both findings in Section 9.
@@ 8.1.1 Ablation: Lying Evaluators and Lying Generators
Table 2: Results for the known (test score if the labels were known) and unknown (output of the No-Data Algorithm) cases. The known case is a reference that in practice does not exist. For the unknown case, if we used the base generator, the performance is exactly the same as in the known case. The flip $\phi$ is set near the accuracy in the known case ($0.6$ for DT; $0.9$ for LLM). This is because the generator acts as an oracle, convincing the verifier every time, and hence the No-Data algorithm reduces to simple evaluation (the known case). However, the success rate is extremely low in DT OOP (unknown): even though there is no way for us to know whether the evaluator’s score is trustworthy, the success rate indicates deception.

### v3

@@ 5 Characterisation
The implication of Theorem 1 is that a good evaluator with relatively high accuracy and success rate, and a well-calibrated $\phi$, will have a true accuracy higher than what the No-Data Algorithm shows. This means that reliable evaluators, as per the algorithm, may be used with the expectation that their ‘true’ performance will be better. Another interpretation of this result is that there is no way to recover unknowable labels when the evaluator is unable to convince the verifier.
It is worth mentioning that the assumption for Theorem 1 around always lying when mislabelling (and hence always being wrong) are not necessarily applicable to all evaluators, since a particularly tricky evaluator could still guess the right label and yet not pass the challenges. We consider this out of scope, since the ability to guess a label without explaining how is not a viable way to establish trust, especially in scenarios without any references.
@@ 6 Experiments
We present two experiments: an empirical test of the theory (Section 6.1), and an application to a realistic scenario (Section 6.2). Ablation experiments are in Appendix D. We used two corpora: the data it knows how to label (in-phenomenon, or IP), and one it doesn’t (out-of-phenomenon; OOP). We use this naming to highlight the fact that the entire measurement is altered by the change in data–and, that, semantically, we are modelling different phenomena (i.e., distinct $f\colon X\rightarrow Y$). We refer to the hypothetical setup where we know the labels of a dataset as knowable (r. unknowable). Knowable scenarios are only to illustrate how the algorithm’s labelling adheres to the ‘true’ predictions, and in practice they would be unavailable. We run the No-Data Algorithm with $r=3$. See Appendix F for prompts and specifics on the models and methodology used.
6.1 Experiment 1: Synthetic Data without Natural Language
In this experiment we aimed to understand the behaviour of the No-Data Algorithm with a corpus close to the theoretical setup. We created two disjoint synthetic datasets over binary strings. This had the advantage of ablating out any potential memorisation concerns, and allowed us to control the rubrics, as shown in Rubric 1. For both OOP and IP, $\sigma$ is the majority vote. The test sets for both IP and OOP included 498 entries and are balanced.
6.1.1 Setup: Evaluators and Verifiers
Evaluators:
We used a decision tree (DT), and the highest-performing LLM we tested (o3-mini; OpenAI 2025a). [^2] The full results, in Appendix D, further include another closed- and two open-source models. Both observed a balanced training split of the IP dataset (DT was trained in it; the LLM observed it with five exemplars). We trained the DT to ensure that the evaluator learnt both IP’s $\sigma$ and $C$, even though during generation it only had access to $C$, and not $\sigma$. As ‘training’, the LLM observed the rubric in the prompt, and was made aware of the aggregation function only for the labelling step. For generation it was asked to pick the best candidate from the (training) dataset instead of outputting a new point, as we observed marked performance decreases otherwise (Appendix D). During OOP, the LLM was given the rubric for IP and an OOP datapoint. We tuned $\phi$ based on the knowable scenarios, to $\phi=0.4,0.1$ for the DT and the LLM, respectively.
Verifier:
We used a rule-matching algorithm that has access to the relevant rubrics (e.g., when testing IP, it observes the rubric for IP), implemented as described in Section 4.1.
@@ 6.1.1 Setup: Evaluators and Verifiers
$c_{0}$ $y=1$ if $x$ has an even number of ones; else 0
$c_{1}$ $y=1$ if $x$ starts with a zero or contains $10101$, but not both; else 0
$c_{2}$ $y=1$ if $x$ has strictly more than five ones; else 0
@@
$c_{0}$ $y=1$ if $x$ contains the substring $111$; else 0
$c_{1}$ $y=1$ if $x$ ends with a one; else 0
$c_{2}$ $y=1$ if $x$ contains the substring $110001$; else 0
Table 1: Rubrics and criteria for our first experiment. Note that $c_{1}$ for the IP rubric is an exclusive-or operation: while this rubric contains three criteria, the total rubric contains five: $\{c_{0},c_{2},c_{1,a},c_{1,b},c_{1}\}$; where $c_{1,a},c_{1,b}$ are the clauses for $c_{1}$, and $c_{1}$ is the exclusive-or itself.
6.1.2 Results
The results are in Table 1. The DT and the LLM adjusted to the predictions: accuracies with the No-Data Algorithm were within 2-2 and 0% of the knowable case in both IP and OOP. The success rate differences between IP (100% for DT; 81% for the LLM) and OOP (r. 5% and 28%) provided better information as to which setup was the one the evaluator truly knew. Of note, which we will cover in Section 7, is the number of flips performed by the algorithm in the DT OOP dataset: 46%.
We also performed ablation studies on the relationship of lying with evaluators and generators; the need for flipping the labels; and the effectiveness of generation strategies in various LLMs. Full details for all studies are available in Appendix D. We found that not flipping the labels led to lower accuracies in the No-Data Algorithm, especially when the evaluator lied and within IP. When the evaluator lied about knowing $\sigma$ or $f$, the success rates were lower in the IP case, but remained steady in the OOP case. The final study showed that LLM generation under this problem led to poor results in comparison to picking a string from a dataset.
@@ 6.1.2 Results
Table 1: Results for the known (test score if the labels were known) and unknown (output of the No-Data Algorithm) cases. The known case is a reference that in practice does not exist. For the unknown case, if we used the base generator, the performance is exactly the same as in the known case. The flip $\phi$ is set near the error rate in the known case ($0.4$ for the DT; $0.1$ for the LLM), since it is arguably the worst value it can have. This is needed in this experiment because the generator acts as an oracle, convincing the verifier every time, and hence the No-Data algorithm reduces to simple evaluation (the known case). However, the success rate is extremely low in DT OOP (unknown): even though there is no way for us to know whether the evaluator’s score is trustworthy, the success rate indicates deception.
6.2 Experiment 2: Low-Resource Language Labelling
6.2.1 Data
We created a multi-domain dataset (1,015 entries total) with randomly-drawn samples from OpenOrca (Lian et al., 2023), MMLU (Hendrycks et al., 2021), OpenCode Ahmad et al. (2025), and WildChat (Zhao et al., 2024). It was then professionally translated and annotated into West Frisian by four native speakers. See Section 9 and Appendix F for details on the corpus creation. West Frisian is a language that, albeit spoken by half a million people, is within the class of ‘exceptionally limited resources’ with ‘virtually no labelled data to use’, as per the taxonomy from Joshi et al. (2020).
Every entry in the corpus is a pair of prompts and outputs (LLM responses from GPT-4o), and the task was to judge whether the output is correct based on a rubric. The rubric, written in natural language, for IP consisted of six commonly-evaluated criteria ranging from ‘the response must be in West Frisian’ to ‘the response must be correct and without syntax errors’. For OOP we wrote a separate, almost-orthogonal rubric (e.g. ‘the response must be in Dutch’, as the model often responded in Dutch instead), and included a compound entry (‘one of B or C must be correct, otherwise the response is zero’), bringing the total criteria to seven. Full rubrics are in Appendix E. The $\sigma$ for this corpus was label 1 if all criteria were 1, otherwise 0, and performed the evaluation with the IP labels. Remark that this is equivalent to having a lying evaluator, in line with the setup from Section 6.1.
The evaluator and verifier was GPT-4.1 (OpenAI, 2025b). We split the dataset into train (500) and test (515). Ahead of running the experiments, we calibrated the prompt strategies that worked best under the IP known case for both the final labels and the per-criterion breakdown. While this is unrealistic (in practice users would not have access to this dataset), as before it allows us to have a ground truth from which to compare the No-Data Algorithm’s results.
Unlike the previous experiment, here neither the verifier nor the evaluator knew the label or the criteria values, and instead both created their own label sets. We observed that presenting exemplars to the model led to poor performance in both the IP and OOP cases, so the prompts do not use them.
6.2.2 Results
The results are in Table 2. They adjusted to the predictions, with high success rate and accuracies in the IP setting, and the converse in the OOP case. There were some nuanced differences when compared to the synthetic experiments from Section 6.1. Namely, the difference in performance was higher when $\phi$ was low (up to -3.9% and -2.2% accuracy in IP and OOP, r.). Nonetheless, success rates remained steady regardless of value of $\phi$. Additionally, we performed an ablation study comparing a model not proficient in West Frisian (Qwen 2.5B VL 7B; Qwen Team 2025). We found that in the IP case, where the model was given the correct rubric (i.e., it did not lie, but was wrong), the success rate was still low, hovering at around 35 in IP; and 1.8 for OOP (where it did lie), for all values of $\phi$. The full results are in Appendix D.3.
IP Acc. / F1 OOP Acc. / F1 IP Succ. / Flips OOP Succ. / Flips
Known
Predicting $y$ 76.3 / 80.3 51.5 / 67.2 –/– –/–
Avg. per $c\in\bar{C}$ 89.5 / 94.0 45.3 / 46.5 –/– –/–
Unknown
GPT-4.1
$\phi=0.7$ 72.4 / 76.6 49.3 / 57.8 87.8 / 9. 3 1.4 / 70.3
$\phi=0.3$ 74.4 / 78.1 49.1 / 39.1 86.8 / 4.9 1.2 / 33.6
$\phi=0.1$ 76.9 / 79.8 50.7 / 16.4 86.2 / 1.8 1.2 / 9.1
Table 2: Performance of the LLM on the natural-language version of the No-Data Algorithm. We report the best prompts in the known case, which in turn informed the prompts we used in the unknown case. When using the predictions for the criteria to determine the label, the accuracy of the models is 80.0% (IP), and 50% (OOP).
7 Discussion
7.1 Theory
Theorem 1 states that when the labeller has high accuracy, a low $\phi$ will be able to reproduce the label set. It is then possible to measure the successes to establish a confidence bound on the aggregate labels provided by the evaluator. It, however, has a different interpretation: when the accuracy of the evaluator is zero, there is no way to determine the labelling, regardless of $\phi$–it is completely unknowable. We argue that this makes sense: otherwise, it would be possible to create information out of thin air, and without any prior knowledge of $f$. Hence the primary purpose of the No-Data Algorithm is to establish trustworthiness, not to label data.
We assumed that $Y$ is a binary set, and that $X$ is the set of $k$-bit binary strings. The first assumption can be easily addressed by ‘binarising’ a set $Y$ with arity >2>2 by class-based separation. The second must be handled carefully from a mathematical point of view. For example, to model natural-language problems it would be inaccurate to assume that the $k$-bit encoding for $x\in X$ will retain correctness by solely working with the binary encoding of the strings. Extra information (e.g., semantics captured in a latent space) must be added to $x$, before proceeding as usual. Remark that, as per our experiments, this is not a concern empirically. Further developments of the theory addressing this and other aspects is in Appendix C.
7.2 Experiments
In both experiments, the results adjusted to the theory’s predictions: both the LLM and DT evaluators consistently had accuracies close to the known cases in IP/OOP, but distinct success rates in IP and OOP. This indicated that the No-Data Algorithm is robust to evaluator strategies attempting deception.
The high number of flips in the $\phi=0.5$ DT OOP case (Section 6.1) and $\phi=0.3$ LLM OOP (Section 6.2) support our interpretation of Theorem 1. Since the model’s accuracy is near-random in the OOP dataset–as indicated by the known case–the verifier will almost always catch a lie. Hence, most labels got flipped based on $\phi$–indeed, the exact proportion marked by $\phi$–and no information was gained from the labelling. It is worth noting that the success rate remained steady regardless of the value of $\phi$, even when the flips and accuracies didn’t. This means that it is an interpretable quantity, while $\phi$ is tuneable to attain ‘good’ labelled values.
Finally, it is worth noting that the evaluation of LLMs with LLMs, as per Section 6.2, depends strongly on the LLMs themselves. This is not a circular argument per se. The judge (of judges) is tuned on the IP case, and asked to evaluate an unkown model in the OOP case, as done in Section 6.2 and Appendix D.3, even when it had near-random performance in this task. This is because they are solving different tasks, and thus the accuracies did adjust to the known cases. From the perspective of future research directions, it does raise the question on what are the results measuring. We come back to this point in the next section.
8 Conclusion
We introduced an algorithm to determine trustworthiness of an evaluator in the case where labelled data does not exist. It works by performing randomly-selected checks evaluating specific parts of the datapoint to be labelled. We showed that these checks were sufficient and necessary to determine (w.h.p.) whether the evaluator knew the labels. Experimentally, the results adjusted to the theoretical predictions, with the No-Data Algorithm being able to determine lying evaluators under various scenarios, both synthetic and realistic. Of note was when the evaluator was randomly guessing the labels: while its accuracy was deceptively close to that of its known case, the success rate of the No-Data Algorithm was low, thus flagging that said labelling could not be trusted. Since this quantity remained steady across all values of $\phi$, it makes it a reliable metric to determine trust without needing tuning.
One implication of the No-Data Algorithm’s correctness proof is that, when the evaluator cannot be fully trusted, the data cannot be annotated. This scenario is unsolvable: there is no prior information on the data, and hence it would be effectively creating information out of thin air otherwise. This is why the algorithm’s focus is on the successes: if an evaluator has low accuracy, but high success rate, it is very likely it is not lying; just wrong (e.g., the IP case in the experiments). Alternatively, high accuracy but low success rate indicates deception (r. the OOP case). Hence, the primary purpose of this algorithm is to establish trustworthiness and is up to the user to decide (to its own margin of trust) whether to accept or reject the labelling.
It is well-known that LLMs are sensitive to prompting. Establishing trust in an LLM evaluator in our setup was equivalent to establishing trust in the prompt, not the model. While effective in the two experiments we ran, this nuance could require extensions to the No-Data Algorithm to determine (w.h.p.) whether it is the model or the prompt what is failing to convince the verifier. As a starting point, we note that our realistic experiment slightly departed from the theoretical work by not assuming that the criteria was known, and instead relied on the evaluator and verifier to generate their own judgement of both $C$ and $\tilde{y}$. Nonetheless, our work has shown that it is possible, theoretically and empirically, to judge an evaluator–such as an LLM-as-a-judge–up to formal (mathematical) guarantees of correctness. The other implication of our work is that the algorithm’s setup allows the verifier to ‘self-calibrate’, by generating its own datapoints, since it knows which challenges will appear. This has both theoretical and practical implications that we leave for future exploration.
9 Ethics
This work’s contributions involve the ability to establish trust in an LLM. We are unaware of any potential misuse of this algorithm, albeit we admit we are unable to cover all potential applications. We, however, argue that releasing the code publicly has more benefits than potential harms. For the data annotation, all annotators were contracted through an annotator services company, and compensated at a rate of 28.7 USD/hr. Annotators were encouraged to take breaks and emphasised quality over speed through the annotation instructions.
10 Reproducibility Statement
All code has been included as part of the supplementary materials. It will be released under the MIT licence. The data will be released on the three licences from the original works. Detailed methodology, included model versioning, is in Appendix F. Prompts are in Appendix F.1, and also in the supplementary materials. Whenever available, we set the temperature to zero to ensure further reproducibility. Given the tendency of closed-source models to update and phase out existing models, we also include experiments with open models in Appendix D. The theoretical work includes the assumptions in Sections 3 and 5, and detailed proofs in Appendix A.
Appendix A Detailed Proofs
A.1 Correctness of the EV protocol
A.1.1 Part 1 (Correctness)
We prove the following lemma:
Lemma 2.
Proof.
(Sufficient) It follows directly from the earlier statement. Suppose there exists a pair $x,x^{\prime}$ such that either (or both) $\not\equiv x^{\prime}, x≇x$$\not\cong x^{\prime}$ and also $f(x)=f(x^{\prime})$. This would imply that (1) either the string , a contradiction.
This concludes the proof. ∎
A.1.2 Part 2 (Robustness)
We prove the following lemma:
Lemma 3.
Proof.
In here we have two scenarios: either the evaluator knows $f$, or it doesn’t, and hence it is lying or cheating.
A.2 Lie 1: Generate a ‘good’ $x^{\prime}$
A.3 Lie 2: $f(x^{\prime})\neq y^{\prime}$
It follows from Lemma 2 that we know that both challenges together are sufficient to ascertain the truthfulness of a label; and from Lemma 3 we know that the ability of the verifier to do this is bounded. Putting both lemmas together we obtain the proof for Lemma 1.
A.4 Correctness Bounds of the No-Data Algorithm
For this, it is easier to work with errors $\epsilon=1-\alpha$, and remark that, from Lemma 1 the first term (undetected, no flip) is given by
This yields:
$\bigcup\Pr[\text{wrong}]$ $\leq\Pr[\text{undetected}]+\Pr[\text{detected, no flip}]$ (7)
$\mathrm{E}[\text{wrong}]$ $\leq\epsilon\left(1-\phi+\phi\left(\frac{1}{4}\right)^{r}\right)\text{, by linearity of expectation.}$ (9)
Appendix B Characterisation Bounds
Lemma 4.
[…]
Proof.
Straightforward. ∎
Appendix C Extensions
C.1 Distance-based Constraints
C.2 $k$-ary Label Sets
To convert a dataset of arity $k>2$, it is necessary to rerun the No-Data algorithm k1k-1 times. On each run, the label set $\{l_{1},\dots,l_{k}\}$ is mapped as $\{l_{1}\},\{l_{2},\dots,l_{k}\}\mapsto\{0\},\{1\}$. In these cases, the runtime of the algorithm is increased by a factor of k1k-1.
Appendix D Full Results
D.1 Ablation: Lying Evaluators and Lying Generators
  1. 1.
  2. 2.
  3. 3.
    approximately know $f$ up to a probability $p=1/10$ (r. it could pass both challenges with probability 1p1-p).
As before, the results indicate that lying evaluators tend to score low in the success rate, even when their accuracies are close to the known case. Remark that (a) the evaluator that lies on knowing $\sigma$ had a success rate much higher than that of lying of knowing $f$; and (b) the evaluator with approximately knowledge of $f$ had a comparatively low (44%) success rate.
This success rate depends on the cardinality of $\bar{C}$: when it is equal to $\mathcal{C}$, the model will always pass Challenge 1 (and fail Challenge 2). However, when $\lvert\bar{C}\rvert>\lvert C\rvert$, there will be criteria that the evaluator cannot possibly solve without knowing how $C(x)$ is constructed based on its internal structure, and hence it will fail them with some probability.
D.2 Ablation: LLM Generation Strategies
Given the LLMs’ ability to follow instructions, it is likely that their performance could be sensitive to how the datapoints are generated during the EV protocol. Hence in this study we compared the performance of both the LLM whose results we presented in the main results (o3-mini) with another LLM, GPT-4o [OpenAI, 2024]. We tested two different generation strategies: picking a matching datapoint from the data, as in Section 6.1.2, and actually generating a datapoint from the rubric. The results are in Table 3. We found that, again, the experimental results adjusted to the predictions, with lying and low-performing models being easily spotted by the No-Data Algorithm. However, we also found that actual generation led to bad performance in both LLMs. We note that this is expected: in general, classifying an element in a language is much harder than generating it [Kleinberg and Mullainathan, 2024, Kalavasis et al., 2025].
o3-mini (IP) o3-mini (OOP) GPT-4o (IP) GPT-4o (OOP)
Picking
Accuracy/F1 97.6 / 97.6 59.0 / 70.0 58.0 / 67.1 56.4 / 65.4
Successes/Flips 81.3 / 1.8 27.9 / 6.0 28.7 / 71.3 10.8 / 8.8
Generating
Accuracy/F1 94.2 / 94.2 60.6 / 65.5 56.6 / 63.8 53.4 / 63.6
Successes/Flips 55.4 / 5.0 28.7 / 7.8 5.8 / 10.8 34.5 / 4.6
Table 3: Ablation study on LLM generation strategies, comparing a high-performing model (o3-mini) with a lower-performing model (GPT-4o) as a LLM-based baseline. We report the highest-performing run at $\phi=0.1$. Both models were able to broadly adjust to their original accuracy, but success rates are markedly different when the models are generating a datapoint versus picking one.
We also performed a study on the overall performance of various models under the generation strategy, comparing them to the known case. This study includes results with DeepSeek [DeepSeek-AI et al., 2025], GPT-4o, and Qwen 2.5 VL 7B [Qwen Team, 2025] (Table 4) Outside of o3-mini, all models–open or not–struggled with this problem, even in the known case. Consequentially, their performance in the unknown case also adjusted to the predictions. When the accuracy was high in OOP settings, the success rate was low–as expected.
o3-mini DeepSeek GPT-4o Qwen
Known Case
IP
[…]
Successes/Flips –/– –/– –/– –/–
OOP
[…]
Successes/Flips –/– –/– –/– –/–
Unknown Case
($\phi=0$)
IP
[…]
Successes/Flips 53.8 / 46.2 12.9 / 87.1 6.6 / 93.4 13.3 / 87
OOP
[…]
Successes/Flips 26.3 / 73.7 16.5 / 39.5 34.1 / 65.9 16.3 / 83.7
($\phi=0.5$)
IP
[…]
Successes/Flips 55.4 / 5.0 12.5 / 44.2 6.6 / 47.8 13.1 / 47.8
OOP
[…]
Successes/Flips 28.7 / 59.3 19.5 / 37.8 33.7 / 34.1 16.3 / 47.8
($\phi=0.9$)
IP
[…]
Successes/Flips 55.4 / 5.0 11.9 / 9.4 5.8 / 10.8 13.1 / 8.8
OOP
[…]
Successes/Flips 28.7 / 7.8 18.1 / 7.6 34.5 / 4.6 16.3 / 8.0
Table 4: Comparison of the No-Data algorithm with respect to various LLMs and values of $\phi$. The known case reflects the LLM’s expected performance in the dataset, while the unknown case shows what the No-Data Algorithm outputs. In the unknown case, all models broadly adjusted to their known case. The models also showed on average low accuracies across the board. However, success rates were markedly different between the known and unknown cases under high accuracy, indicating possible deception. When the accuracy was low, success rates remained low regardless of the setting. For all of the above, we used the generating strategy.